OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (6, -4) and (18, 10)?

OpenStudy (anonymous):

what is the midpoint between those 2 points?

OpenStudy (rmrjr22):

(x-a)^2 + (y-b)^2 = r^2

OpenStudy (rmrjr22):

\[(x2+x1)/2, (y2+y1)/2 = midpoints\]

OpenStudy (anonymous):

12,14?

OpenStudy (rmrjr22):

yes

OpenStudy (rmrjr22):

no.... the second one is wrong

OpenStudy (anonymous):

7?

OpenStudy (rmrjr22):

10 + (-4) = ?

OpenStudy (anonymous):

6

OpenStudy (rmrjr22):

yup

OpenStudy (rmrjr22):

best response? ;)